# Quantitative Certification of Bias in Large Language Models

### v2

@@ Abstract
Warning: This paper contains model outputs that are offensive in nature. Large Language Models (LLMs) can produce biased responses that can cause representational harms. However, conventional studies are insufficient to thoroughly evaluate LLM bias, as they can not scale to large number of inputs and provide no guarantees. Therefore, we propose the first framework, QuaCer-B (Quantitative Certification of Bias) that certifies LLMs for bias on distributions of prompts. A certificate consists of high-confidence bounds on the probability of unbiased LLM responses for any set of prompts mentioning various demographic groups, sampled from a distribution. We illustrate the bias certification for distributions of prompts created by applying varying prefixes drawn from a prefix distributions, to a given set of prompts. We consider prefix distributions for random token sequences, mixtures of manual jailbreaks, and jailbreaks in the LLM’s embedding space to certify bias. We obtain non-trivial certified bounds on the probability of unbiased responses of SOTA LLMs, exposing their vulnerabilities over distributions of prompts generated from computationally inexpensive distributions of prefixes.
@@ 1 Introduction
Text-generating Large Language Models (LLMs) are recently being deployed in user-facing applications, such as chatbots (Lee et al., 2023). LLM-powered chatbots, like ChatGPT (Brown et al., 2020a) and Perplexity AI (Perplexity, 2023), are popular for their ability to produce human-like texts (Shahriar and Hayawi, 2023). The underlying LLMs are safety-trained (Wang et al., 2023) to avoid generating harmful content. However, despite safety training, they have been shown to produce texts that exhibit social biases and stereotypes (Kotek et al., 2023; Manvi et al., 2024; Hofmann et al., 2024). Such texts can result in representational harms (Suresh and Guttag, 2021; Blodgett et al., 2020) to protected demographic groups (a subset of the population that is negatively affected by bias). Representational harms include stereotyping, denigration, and misrepresentation of historically and structurally oppressed demographic groups. Although “representational harms are harmful in their own right” (Blodgett et al., 2020), as they can socially impact individuals and redefine social hierarchies, the resulting allocation harms can lead to economic losses to the protected groups and are hence regulated by anti-discrimination laws such as (Sherry, 1965). Language is considered an important factor for labeling, modifying, and transmitting beliefs about demographic groups and can seldom result in new social hierarchies and amplify representational harm. Hence, with the rising popularity of LLMs, it is important to formally evaluate their biases to effectively mitigate representational harms due to them.
@@
Prior work has primarily focused on benchmarking the performance of LLMs (Liang et al., 2023; Wang et al., 2024; Mazeika et al., 2024) and adversarial attack generation (Sheng et al., 2020; Zou et al., 2023; Vega et al., 2023; Wallace et al., 2019). While these methods provide some empirical insights into LLM bias, they have several fundamental limitations (McIntosh et al., 2024; Yang et al., 2023) such as — (1) Limited test cases: Benchmarking consists of evaluating several but limited number of test cases. Due to its enumerative nature, benchmarking can not scale to the prohibitively large sets of prompts that can elicit bias from LLMs. Adversarial attacks, on the other hand, identify only a few worst-case examples of bias, which do not inform about the overall biases from large input sets; (2) Test set leakage: LLMs may have been trained on the popular benchmarking datasets, thus resulting in biased evaluation; (3) Lack of guarantees. Benchmarking involves empirical estimation that does not provide any formal guarantees of generalization over any input sets. Similarly, adversarial attacks give limited insights as they can show existence of problematic behaviors on individual inputs but do not quantify the risk of obtaining biased LLM responses.
@@
This work. In this work, we propose an alternative to benchmarking and adversarial attacks — certifying LLMs for bias which gives formal guarantees. Certification operates on a prohibitively large set of inputs, represented as a specification. As specifications define inputs mathematically through operators over the vocabulary of LLMs, certification can provide guarantees on the behavior of the target model that generalize to unseen inputs satisfying the specification. With guarantees, we can be better informed about the available models before deploying them in public-facing applications.
@@
Key challenges. (1) There are no existing precise mathematical representations of large sets of prompts to make practical specifications. (2) State-of-the-art neural network certifiers (Wang et al., 2021; Singh et al., 2019) currently do not scale to LLMs as they require white-box access to the model and lose precision significantly for larger models, resulting in inconclusive results.
@@
Our approach. Given the diversity of prompts handled by LLMs, there will always be a few cases where the LLM output will be biased (e.g., these can be found with adversarial attacks (Zou et al., 2023; Chao et al., 2023)). Hence, we believe that LLM certification must be quantitative (Li et al., 2022a; Baluta et al., 2021) and focus on the following question:
What is the probability of LLM responses being unbiased for any random set of prompts each mentioning different demographic groups?
@@
Figure 1: (Overview of QuaCer-B): QuaCer-B is a quantitative certification framework to certify the bias in the responses of a target LLM for a random set of prompts that differ by their sensitive attribute. In specific instantiations, QuaCer-B samples a (a) set of prefixes from a given distribution and prepends them to a prompt set to form (b) the prompts given to the target LLM. (c) The target LLM’s responses are checked for bias by a bias detector, (d) whose results are fed into a certifier. (e) The certifier computes bounds on the probability of obtaining biased responses from the target LLM for any set of prompts formed with a random prefix from the distribution.
@@
We study bias over a set of prompts with varying demographic groups, inspired from Counterfactual Fairness (Kusner et al., 2018). In particular, we identify the biases across the LLM responses for counterfactual prompts (that differ only by the mentioned demographic groups) to capture the difference in their outputs caused by just varying demographic groups in the prompts. Exactly computing the probability of unbiased responses is infeasible due to the large number of possible sets of counterfactual prompts over which the biased behavior has to be determined. One can try to compute deterministic lower and upper bounds on the probability (Berrada et al., 2021). However, this is expensive and requires white-box access making it not applicable to popular, SOTA but closed-source LLMs such as GPT-4 (Achiam et al., 2023). Therefore, we focus on black-box probabilistic certification that estimates the probability of unbiased responses with high confidence bounds. We develop the first general specification and certification framework, QuaCer-B [^1] Quantitative Certification of Bias for LLM bias, applicable to both open and closed-source LLMs. Figure 1 gives an overview of our framework. Our specifications encode the probability of unbiased LLM responses for counterfactual prompts. We demonstrate QuaCer-B with $3$ exemplar kinds of specifications, each characterized by a specific kind of distribution of prefixes applied to a fixed set of counterfactual prompts. The distributions of prefixes we present are — random sequence of tokens, mixture of popular jailbreaks, and jailbreak perturbations in embedding space. The first two are model-agnostic specifications and hence apply to both open and closed-source models. However, the third one requires access to the embeddings and the ability to prompt LLMs with embeddings, and hence mostly applies to only open-source models. The prefixes are described further in Section 3.1.
@@
We certify the proposed specifications leveraging confidence intervals. Our certifier samples several counterfactual prompt sets from the distribution given in the specification and generates high-confidence bounds on the probability of getting unbiased responses from the target LLM for any random prompt set in the distribution.
@@
  • We design novel specifications that quantify the desirable property of low bias in the responses of LLMs with varying prefixes. The prefixes to given counterfactual prompts can be drawn from $3$ distributions — (1) random, (2) mixture of jailbreaks, and (3) jailbreak perturbations in the embedding space.
  • We develop a probabilistic black-box certifier QuaCer-B, applicable to both open and closed-source models, for quantifying bias in LLM output. QuaCer-B leverages confidence intervals (Clopper and Pearson, 1934) to generate high-confidence bounds on the probability of obtaining unbiased responses from the target LLM, given any random set of prompts from the distribution given in the specification.
  • We use QuaCer-B to reveal novel insights into the biases of popular LLMs, for different specifications (Section 5). We find that the safety alignment of SOTA models is easily circumvented with several prefixes in the distributions given in our specifications, especially those involving mixture of jailbreaks and jailbreak perturbations in the embedding space. These distributions are inexpensive to sample from, but can effectively bring out biased behaviors from SOTA models. This shows the existence of simple, bias-provoking distributions for which no defenses exist currently. We provide quantitative measures for the fairness (lack of bias) of SOTA LLMs, which hold with high confidence. We find that there are no consistent trends in the fairness of models with the scaling of their sizes, hence suggesting that the quality of alignment techniques could be a more important factor than size for fairness. Our implementation is available at https://github.com/uiuc-focal-lab/QuaCer-B.
@@ 2 Background
2.1 Large Language Models
Large Language Models (LLMs) are autoregressive foundation models that perform next-token prediction. That is, given a sequence of tokens $t_{1},\dots,t_{k}$, they give a distribution over the vocabulary for the next token, $P[t_{k+1}\mid t_{1},\dots,t_{k}]$. They are typically fine-tuned for instruction-following capabilities to be used such as chatbots. These instruction-finetuned models are further aligned with human feedback (Wang et al., 2023; Ouyang et al., 2022) to make their responses safe and unbiased. We certify these instruction-tuned (Zhang et al., 2024), aligned models for fairness, as they are typically used in public-facing applications.
@@ 2.2 Clopper-Pearson confidence intervals
Clopper-Pearson confidence intervals (Clopper and Pearson, 1934) provide lower and upper bounds on the probability of success parameter of a Bernoulli random variable with probabilistic guarantees. The bounds are obtained with $n$ independent and identically distributed observations of the random variable, in which $k(\leq n)$ successes are observed. The confidence interval is constructed to bound the value of $p$ with $[\hat{p_{l}},\hat{p_{u}}]$ with at least a given confidence level $(1-\gamma)100\%$. That is, $Pr\{p\in[\hat{p_{l}},\hat{p_{u}}]\}\geq(1-\gamma)$. $\gamma$ is the error probability by which the true value of $p\notin[\hat{p_{l}},\hat{p_{u}}]$. These confidence intervals are obtained by binomial statistical hypothesis testing for the value of $p$, where the lowest and highest values are $\hat{p_{l}}$ and $\hat{p_{u}}$ respectively, with the given confidence $1-\gamma$.
@@ 3 Formalizing bias certification
We develop a general framework, QuaCer-B, to specify and quantitatively certify the bias in the text generated by (Large) Language Models. QuaCer-B formalizes bias with specifications — precise mathematical representations that define the desirable property (absence of bias) in large sets of inputs. Bias is defined with respect to demographic groups that are subsets of the human population that share an identity trait, that could be biological, contextual, or socially constructed (Gallegos et al., 2024b). Bias consists of disparate treatment or outcomes when varying the demographic groups in the inputs to the target model (Gallegos et al., 2024b; Sheng et al., 2019; Smith et al., 2022). In the context of autoregressive LMs, we consider text generation bias consisting of stereotyping, misrepresentation, derogatory language, etc. This can primarily result in representational harms, which can subsequently cause allocation harms (Gallegos et al., 2024b). To apply certification to closed-source LMs as well, we study extrinsic bias (Cao et al., 2022) that manifests in the final generated responses of the models.
@@ 3.1 Bias specification
Next, we formally specify the lack of bias in the outputs of language models (LMs). In particular, we want the text generated by LMs to not exhibit any semantic disparities caused by mentioning different demographic groups in the prompts. Hence, our bias specification is motivated by Counterfactual Fairness (Kusner et al., 2018). Consider a given identity trait $\mathcal{I}$ such as gender, race, etc. (that are often the basis of social bias). $\mathcal{I}$ categorizes the human population into $m$ subsets called demographic groups $\mathcal{G}_{1},\dots,\mathcal{G}_{m}$, each differing by the value of the identity trait. Each demographic group $\mathcal{G}$ is characterized/recognized by several synonymous strings, called sensitive attributes $\mathcal{G}^{\mathcal{A}}$ (Li et al., 2024). For example, the sensitive attributes for the demographic group corresponding to the female gender are woman, female, etc. We select any one sensitive attribute of a demographic group $\mathcal{G}$ to represent $\mathcal{G}$. Let the resulting set of sensitive attributes, each corresponding to a demographic group, be $\mathcal{A}=\{\mathcal{A}_{1},\dots,\mathcal{A}_{m}\}$, where $\mathcal{A}_{i}\in\mathcal{G}^{\mathcal{A}}_{i}$. Our specifications are for counterfactual inputs (prompts) (Gallegos et al., 2024b) that differ only by the sensitive attributes in them.
Let $\mathcal{L}$ be the target LM and $\mathcal{V}$ be its vocabulary. Consider a set of prompts $\mathcal{P}=\{\mathcal{P}_{1},\dots\mathcal{P}_{s}\}$, $s>1,\mathcal{P}\subset\mathcal{V}^{[1,c]}$, where $c$ is the context length of $\mathcal{L}$ and $\mathcal{V}^{[1,c]}$ is a sequence of elements of $\mathcal{V}$ having length $\in[1,c]$. Let each prompt in $\mathcal{P}$ contain a unique sensitive attribute from $\mathcal{A}$ such that overall $\mathcal{P}$ represent more than $1$ distinct demographic groups represented in $\mathcal{A}$. Let each prompt be $\mathcal{P}_{i}=X_{i}\cup\mathcal{A}_{i}$, where $X_{i}$ is the part of the prompt that is independent of sensitive attributes and $\mathcal{A}_{i}$ consists of a sensitive attribute. We consider only prompts that can be decomposed into parts with and without sensitive attributes respectively. To generalize to closed-source LMs, we assume $\mathcal{L}$ to be a black-box system that can only be queried, when specifying and certifying bias. The black-box assumption renders any symbolic analysis (Mirman et al., 2020) infeasible and hence allows only for analysis with input-output examples.
@@
(Counterfactual prompt set). A set of prompts $\mathcal{P}=\{\mathcal{P}_{1}=X_{1}\cup\mathcal{A}_{1},\dots,\mathcal{P}_{s}=X_{s}\cup\mathcal{A}_{s}\}$ is called counterfactual when: (1) $\forall i,j\in[1,s]\ldotp X_{i}=X_{j}$; (2) $\exists i\in[1,s]\ldotp\forall j\in([1,s]\setminus i)\ldotp\mathcal{A}_{i}\neq\mathcal{A}_{j}$; (3) For an unbiased text generator $f$, $\forall i\in[1,s]\ldotp f(\mathcal{P}_{i})=f(X_{i})$. That is, the prompts only differ in the sensitive attributes that are ideally unrelated to the overall query and the set of prompts $\mathcal{P}$ represent more than sensitive attribute from $\mathcal{A}$.
We specify bias over counterfactual prompt sets (Definition 3.1), which can be prohibitive in number for typical context lengths. This is because, the common part of counterfactual prompt sets, $X$ can be any element from $\mathcal{V}^{[1,c]}$ which contains $\approx 10^{10000}$ elements for $c=2k$. Hence, enumerative specifications (which specify the desired behavior on all inputs) are impractical, as they cannot be scalably certified without symbolic analysis for large number of inputs. Hence, we define probabilistic specifications for the probability of unbiased responses from $\mathcal{L}$, for which we provide a certification algorithm in Section 3.2. Let $\Delta$ be a sampleable probability distribution over $\wp(\mathcal{V}^{[1,c]})$ (power set of prompts) having non-zero support on some counterfactual prompt sets $\mathcal{P}$. We define probabilistic specifications for bias in $\mathcal{L}$ over $\Delta$. The specification is agnostic to $\Delta$’s sampler, as long as it generates independent and identically distributed samples. We provide examples of $\Delta$ in Section 4.
Let $\mathcal{D}$ be a user-defined bias detection function that can identify stereotypes/disparity in given texts for different sensitive attributes in $\mathcal{A}$. Let $\mathcal{D}$ evaluate to zero for unbiased inputs (appropriate scaling and shifting of $\mathcal{D}$ can be done to satisfy this criterion). We leave $\mathcal{D}$ as a parameter of the specification, as different domains can have varying notions of bias and the stakeholders can decide which notion is most suitable for their usecase (Anthis et al., 2024).
Overall, we give our quantitative specification as the probability of unbiased responses (as measured by $\mathcal{D}$) when $\mathcal{L}$ is independently prompted with each element of a randomly sampled counterfactual prompts set from $\Delta$ (1). The certificate $\mathcal{C}$ is defined for the user-defined parameters $\Delta$, $\mathcal{D}$, and $\mathcal{L}$.

### v3

# Certifying Counterfactual Bias in LLMs

@@ Abstract
Warning: This paper contains model outputs that are offensive in nature. Large Language Models (LLMs) can produce biased responses that can cause representational harms. However, conventional studies are insufficient to thoroughly evaluate biases across LLM responses for different demographic groups (a.k.a. counterfactual bias), as they do not scale to large number of inputs and do not provide guarantees. Therefore, we propose the first framework, LLMCert-B that certifies LLMs for counterfactual bias on distributions of prompts. A certificate consists of high-confidence bounds on the probability of unbiased LLM responses for any set of counterfactual prompts — prompts differing by demographic groups, sampled from a distribution. We illustrate counterfactual bias certification for distributions of counterfactual prompts created by applying prefixes sampled from prefix distributions, to a given set of prompts. We consider prefix distributions consisting random token sequences, mixtures of manual jailbreaks, and perturbations of jailbreaks in LLM’s embedding space. We generate non-trivial certificates for SOTA LLMs, exposing their vulnerabilities over distributions of prompts generated from computationally inexpensive prefix distributions.
@@ 1 Introduction
Text-generating Large Language Models (LLMs) are recently being deployed in user-facing applications, such as chatbots (Lee et al., 2023; Brown et al., 2020a; Gemini Team, 2024) where they are popular for producing human-like texts (Shahriar and Hayawi, 2023). These LLMs are safety-trained (Wang et al., 2023) to avoid generating harmful content. However, despite safety training, they can produce texts that exhibit social biases and stereotypes (Kotek et al., 2023; Manvi et al., 2024; Hofmann et al., 2024). Such texts can result in representational harms (Suresh and Guttag, 2021; Blodgett et al., 2020) to protected demographic groups (a subset of the population that is negatively affected by bias). Representational harms include stereotyping, denigration, and misrepresentation of historically and structurally oppressed demographic groups. Although “representational harms are harmful in their own right” (Blodgett et al., 2020), as they can socially impact individuals and redefine social hierarchies, the resulting allocation harms (Gallegos et al., 2024a) can lead to economic losses to protected groups and are therefore regulated by anti-discrimination laws such as (Sherry, 1965). Language is considered an important factor for labeling, modifying, and transmitting beliefs about demographic groups and can result in reinforcement of social inequalities (Rosa and Flores, 2017). Hence, with rising popularity of LLMs, it is important to formally evaluate their biases to effectively mitigate representational harms resulting from them (Lee et al., 2024). We focus on counterfactual bias, inspired from Kusner et al. (2018), which assesses semantic differences across LLM responses caused by varying demographic groups mentioned in prompts (counterfactual prompts).
@@
Prior work has focused on benchmarking LLM performance (Liang et al., 2023; Wang et al., 2024; Mazeika et al., 2024) and adversarial attacks (Sheng et al., 2020; Zou et al., 2023; Vega et al., 2023; Wallace et al., 2019). While these methods provide some empirical insights into LLM bias, they have several fundamental limitations (McIntosh et al., 2024; Yang et al., 2023) such as — (1) Limited test cases: Benchmarking consists of evaluating several but limited number of test cases. Due to its enumerative nature, benchmarking can not scale to prohibitively large numbers of prompts that can elicit bias from LLMs. Adversarial attacks identify only few worst-case examples, which do not inform about the overall biases from large input sets; (2) Test set leakage: LLMs may have been trained on popular benchmarking datasets, thus resulting in incorrect evaluation; (3) Lack of guarantees. Benchmarking involves empirical estimation without any formal guarantees of generalization over any input sets. Similarly, adversarial attacks give limited insights as they can show existence of problematic behaviors on individual inputs but do not quantify the risk of biased LLM responses.
@@
This work. We propose an alternative to benchmarking and adversarial attacks — certifying LLMs for bias, with formal guarantees. Certification operates on a prohibitively large set of inputs, represented succinctly as a specification. Specifications define inputs mathematically through operators over the vocabulary of LLMs. Certification can provide guarantees on the target model’s behavior that generalize to unseen inputs satisfying the specification. With guarantees, we can be better informed about the risks associated with LLMs before deploying them in public-facing applications.
@@
Key challenges. (1) There are no existing precise mathematical representations of large sets of (counterfactual) prompts to make practical specifications. (2) State-of-the-art neural network certifiers (Wang et al., 2021; Singh et al., 2019) currently do not scale to LLMs as they require white-box access to the model and lose precision significantly for larger models, resulting in inconclusive results.
@@
Our approach. Given the diversity of LLM prompts, there will always be some cases where the LLM output will be biased (e.g., found by adversarial attacks (Zou et al., 2023)). Hence, we believe that LLM certification must be quantitative (Li et al., 2022a; Baluta et al., 2021) and study the question:
What is probability of unbiased LLM responses for any counterfactual prompt set?
@@
Figure 1: (Overview of LLMCert-B): LLMCert-B is a quantitative certification framework to certify bias in target LLM’s responses to a random set of prompts differing only by a sensitive attribute. In specific instantiations, LLMCert-B samples a (a) set of prefixes from a given distribution and prepends them to a prompt set to form (b) the prompts given to the target LLM. (c) The target LLM’s responses are checked for bias by a bias detector, (d) whose results are fed into a certifier. (e) Certifier computes bounds using the Clopper-Pearson method (Clopper and Pearson, 1934) on probability of unbiased LLM responses for any set of prompts formed with a random prefix from prefix distribution.
@@
Exactly computing the probability of unbiased responses is infeasible due to the large number of possible counterfactual prompt sets over which the biased behavior has to be determined. One can try to compute deterministic lower and upper bounds on the probability (Berrada et al., 2021). However, this is expensive and requires white-box access making it not applicable to popular, SOTA but closed-source LLMs such as GPT-4 (Achiam et al., 2023). Therefore, we focus on black-box probabilistic certification that estimates the probability of unbiased responses over a given distribution of counterfactual prompt sets with high confidence bounds. We develop the first general specification and certification framework, LLMCert-B [^1] LLM Certification of Bias for counterfactual bias in LLMs, applicable to both open and closed-source LLMs (see Figure 1 for an overview). Our specifications over counterfactual prompt sets are the first relational properties (Barthe et al., 2011) for trustworthy LLMs. We demonstrate LLMCert-B with $3$ kinds of specifications, each with counterfactual prompt set distributions formed by adding prefixes sampled from given distribution of prefixes to a fixed set of counterfactual prompts. We present distributions of prefixes consisting — random token sequences, mixtures of popular jailbreaks, and jailbreak perturbations in embedding space (§ 3.1). The first two are model-agnostic specifications and hence apply to both open and closed-source models. However, the third one requires access to the embeddings and the ability to prompt LLMs with embeddings, and hence applies to only open-source models. The mixture and embedding space jailbreak prefix distributions contain effective, manually designed jailbreaks and their perturbations, which are potential jailbreaks, in their sample space and hence assess LLMs’ biases in adversarial settings.
@@
We certify the target LLM leveraging confidence intervals. LLMCert-B samples several counterfactual prompt sets from the specified prompt distribution and generates high-confidence bounds on the probability of unbiased LLM responses for any random counterfactual prompt set in the distribution.
@@
  • We design novel specifications that quantify the desirable relational property of low counterfactual bias in LLM responses over counterfactual prompts in a specified distribution. We illustrate such specifications with distributions of counterfactual prompt sets constructed with potentially adversarial prefixes. The prefixes are drawn from $3$ distributions — (1) random, (2) mixture of jailbreaks, and (3) jailbreak perturbations in the embedding space.
  • We develop the first probabilistic black-box certifier LLMCert-B, applicable to both open and closed-source models, for quantifying counterfactual bias in LLM responses. LLMCert-B leverages confidence intervals (Clopper and Pearson, 1934) to generate high-confidence bounds on the probability of obtaining unbiased responses from the target LLM, given any random set of counterfactual prompts from the distribution given in the specification.
  • We find that the safety alignment of SOTA LLMs is easily circumvented with several prefixes in the distributions given in our specifications, especially those involving mixture of jailbreaks and jailbreak perturbations in the embedding space (§ 5). These distributions are inexpensive to sample from, but can effectively bring out biased behaviors from SOTA models. This shows the existence of simple, bias-provoking distributions for which no defenses exist currently. We provide quantitative measures for the fairness (lack of bias) of SOTA LLMs, which hold with high confidence. We find that there are no consistent trends in the fairness of models with the scaling of their sizes, hence suggesting that the quality of alignment techniques could be a more important factor than size for fairness. Our implementation is available at https://github.com/uiuc-focal-lab/LLMCert-B and we provide guidelines for using our framework for practitioners in Appendix A.
@@ 2 Background
2.1 Large Language Models (LLMs)
LLMs are autoregressive models for next-token prediction. Given a sequence of tokens $t_{1},\dots,t_{k}$, they give a probability distribution over their vocabulary for the next token, $P[t_{k+1}\mid t_{1},\dots,t_{k}]$. They are typically fine-tuned for instruction-following (Zhang et al., 2024) and aligned with human feedback (Wang et al., 2023; Ouyang et al., 2022) to make their responses safe. We certify instruction-tuned, aligned LLMs for counterfactual bias, as they are typically used in public-facing applications.
@@ 2.2 Clopper-Pearson confidence intervals
Clopper-Pearson confidence intervals (Clopper and Pearson, 1934) provide lower and upper bounds $[p_{l},p_{u}]$ on the probability of success parameter $p$ of a Bernoulli random variable with probabilistic guarantees. The bounds are obtained with $n$ independent and identically distributed observations of the random variable, in which $k(\leq n)$ successes are observed. The confidence interval is such that $Pr\{p\in[[p_{l},p_{u}]\}\geq(1-\gamma)$. $\gamma\in(0,1)$ is the (small) permissible error probability by which the true value of $p\notin[p_{l},p_{u}]$. The confidence intervals are obtained by statistical hypothesis testing for $p$, where the lowest and highest values are $p_{l}$ and $p_{u}$ respectively, with the given confidence $1-\gamma$.
@@ 3 Formalizing bias certification
We develop a general framework, LLMCert-B, to specify and quantitatively certify counterfactual bias in text generated by (Large) Language Models (LMs). LLMCert-B formalizes bias with specifications — precise mathematical representations that define the desirable property (absence of bias) in large sets of inputs. Bias is defined with respect to demographic groups — subsets of human population sharing an identity trait, that could be biological or socially constructed (Gallegos et al., 2024b). Bias consists of disparate treatment or outcomes when varying the demographic groups in inputs to the target LM. For autoregressive LMs, we consider text generation bias consisting of stereotyping, misrepresentation, derogatory language, etc, that can result in representational harms (Gallegos et al., 2024b). To apply certification to closed-source LMs as well, we study extrinsic bias (Cao et al., 2022) that manifests in the final LM response texts. See Appendix I for a detailed discussion on bias in ML.
@@ 3.1 Bias specification
Next, we formally specify the lack of bias in the responses of language models. Unbiased LM responses do not exhibit semantic disparities owing to specific demographic groups in the prompts (Gallegos et al., 2024b; Sheng et al., 2019; Smith et al., 2022). Our bias specification is motivated by Counterfactual Fairness (Kusner et al., 2018). Consider a given identity trait $\mathcal{I}$ such as gender, race, etc. (that are often basis of social bias). $\mathcal{I}$ categorizes the human population into $m$ subsets called demographic groups $\mathcal{G}_{1},\dots,\mathcal{G}_{m}$, each differing by the identity trait. Each demographic group $\mathcal{G}$ is characterized/recognized by several synonymous strings in the society, called sensitive attributes $\mathcal{G}^{\mathcal{A}}$ (Li et al., 2024). For example, the sensitive attributes for the demographic group corresponding to the female gender are woman, female, etc. We select any one sensitive attribute of a demographic group $\mathcal{G}$ to represent $\mathcal{G}$. Let the resulting set of sensitive attributes, each corresponding to a demographic group, be $\mathcal{A}=\{\mathcal{A}_{1},\dots,\mathcal{A}_{m}\}$, where $\mathcal{A}_{i}\in\mathcal{G}^{\mathcal{A}}_{i}$. Our specifications are for counterfactual prompts (Gallegos et al., 2024b) that differ only by the sensitive attributes in them.
Let $\mathcal{V}$ be the vocabulary of the target LM $\mathcal{L}$. Consider a set of prompts $\mathcal{P}=\{\mathcal{P}_{1},\dots\mathcal{P}_{s}\}$, $s>1,\mathcal{P}\subset\mathcal{V}^{c}$, where $c$ is the context length of $\mathcal{L}$ and $\mathcal{V}^{c}$ is a sequence of elements of $\mathcal{V}$ having length $\in[1,c]$. Let each prompt in $\mathcal{P}$ contain a unique sensitive attribute from $\mathcal{A}$ such that overall $\mathcal{P}$ represent more than $1$ distinct demographic groups represented by $\mathcal{A}$. Let each prompt be $\mathcal{P}_{i}=X_{i}\cup\mathcal{A}_{i}$, where $X_{i}$ is the part of $\mathcal{P}_{i}$ independent of sensitive attributes and $\mathcal{A}_{i}$ is a sensitive attribute. We consider only prompts that can be decomposed into parts with and without sensitive attributes respectively. To generalize to closed-source LMs, we assume $\mathcal{L}$ to be a black-box system that can only be queried, when specifying and certifying bias. The black-box assumption renders any symbolic analysis (Mirman et al., 2020) infeasible and hence allows only for analysis with input-output examples.
@@
(Counterfactual prompt set). A set of prompts $\mathcal{P}=\{\mathcal{P}_{1}=X_{1}\cup\mathcal{A}_{1},\dots,\mathcal{P}_{s}=X_{s}\cup\mathcal{A}_{s}\}$ is called counterfactual when: (1) $\forall i,j\in[1,s]\ldotp X_{i}=X_{j}$; (2) $\exists i\in[1,s]\ldotp\forall j\in([1,s]\setminus i)\ldotp\mathcal{A}_{i}\neq\mathcal{A}_{j}$; (3) For an unbiased text generator $f$, $\forall i\in[1,s]\ldotp f(\mathcal{P}_{i})=f(X_{i})$. That is, the prompts only differ in the sensitive attributes that are ideally unrelated to the overall query and the set of prompts $\mathcal{P}$ represent more than $1$ sensitive attributes from $\mathcal{A}$.
We specify bias over counterfactual prompt sets (Definition 3.1). These exclude prompts where sensitive attributes are important to answer the overall query, such as “What steps should I take to prepare for becoming a parent?", as semantically different answers, dependent on the sensitive attribute, are acceptable for such cases. All possible counterfactual prompt sets can be prohibitive in number for typical context lengths. This is because, the common part of counterfactual prompt sets, $X$ can be any element from $\mathcal{V}^{c}$ which contains $\approx 10^{10000}$ elements for $c=2k$. Hence, enumerative specifications (which specify the desired behavior on all inputs) are impractical, as they cannot be scalably certified without symbolic analysis for large number of inputs. Hence, we define probabilistic specifications for the probability of unbiased responses from $\mathcal{L}$, for which we provide a certification algorithm in Section 3.2. Let $\Delta$ be a sampleable discrete probability distribution over $\wp(\mathcal{V}^{c})$ (power set of prompts) having non-zero support on some counterfactual prompt sets $\mathcal{P}$. We define probabilistic specifications for bias in $\mathcal{L}$ over $\Delta$. The specification is agnostic to $\Delta$’s sampler, as long as it generates independent and identically distributed samples. We show examples of $\Delta$ in Section 4.
Let $\mathcal{D}$ be a user-defined bias detection function that can identify stereotypes/disparity in given texts for different sensitive attributes in $\mathcal{A}$. Let $\mathcal{D}$ evaluate to zero for unbiased inputs (scaling and shifting of $\mathcal{D}$ can be done to satisfy this). We leave $\mathcal{D}$ as a parameter of the specification, as different domains can have varying notions of bias and stakeholders can decide the most suitable notion (Anthis et al., 2024). Overall, our quantitative specification is the probability of unbiased responses (measured by $\mathcal{D}$) when $\mathcal{L}$ is independently prompted with each element of a randomly sampled counterfactual prompts set from $\Delta$ (1). Certificate $\mathcal{C}$ is an evaluation/estimation of the specified probability of unbiased responses, along with samples of LLM responses, for user-defined parameters $\Delta$, $\mathcal{D}$, and $\mathcal{L}$.